How much water (density = 1.00 X 103 kg/m3 ) must be displaced to float a cubic block of wood (density 655 kg/m3) that is 15.0 cm on a side?

1
Expert's answer
2021-01-10T18:28:47-0500

Answer

Using buoyancy force equation

"m_wg=m_{wood}g"

Where we indicates for water

"V_w \rho_w=V_{wood} \rho_{wood}"

"V_w =\frac{V_{wood} \rho_{wood}}{\rho_w}"

"=\frac{.15\times.15\times.15\times665}{1000}=0.0022m^3"